REPRESENT THE GIVEN SITUATION IN THE FORM OF QUADRATIC EQUATION-
THE HYPOTENUSE OF A RIGHT TRIANGLE IS 1M LESS THAN TWICE THE SHORTEST SIDE.IF THE THIRD SIDE IS 1M MORE THAN THE SHORTEST SIDE.

Answer by nerdybill(7384)

REPRESENT THE GIVEN SITUATION IN THE FORM OF QUADRATIC EQUATION-
THE HYPOTENUSE OF A RIGHT TRIANGLE IS 1M LESS THAN TWICE THE SHORTEST SIDE.IF THE THIRD SIDE IS 1M MORE THAN THE SHORTEST SIDE.
.
Let x = length of shortest side
then
2x-1 = hypotenuse
x+1 = third side
.
applying Pythagorean theorem:
x^2 + (x+1)^2 = (2x-1)^2
x^2 + (x+1)(x+1) = (2x-1)(2x-1)
x^2 + x^2+2x+1 = 4x^2-4x+1
2x^2+2x+1 = 4x^2-4x+1
2x+1 = 2x^2-4x+1
1 = 2x^2-6x+1
0 = 2x^2-6x
0 = x^2-3x (this is what they're looking for)